# Leveraging multi-source and heterogeneous signals for fatigue detection

### v1

@@ 4.2.1 Regression‑Based Imputation Model Selection
4.2.2 Validity of Regression‑Imputed Data Versus Noise Baselines
The second set of experiments is designed to not only verify Theorems 1 and 2 but also demonstrate that imputed data can effectively support classifier tasks. In this case, VPFD is chosen with its gaze modality masked and reconstructed using three different methods: MLP-predicted values, MLP prediction with Gaussian noise, and pure random noise. We refer to the modified datasets as $\text{VPFD}_{\text{MLP}}$, $\text{VPFD}_{\text{MLP+noise}}$, and $\text{VPFD}_{\text{noise}}$, respectively. The added noise follows a zero‑mean Gaussian distribution whose standard deviation is twice the largest absolute value in the original VPFD gaze data. Finally, we train four separate MLP classifiers using 80% of the original VPFD dataset and each of the three modified versions, respectively. All models are evaluated on the same test set—20% of the original VPFD data—to ensure a consistent basis for comparison. Table 3 clearly shows that the classifier trained on regression-generated data maintains a similar performance to the one trained on the original data. This demonstrates the potential of data synthesis as a solution for addressing missing data. However, the introduction of noise, whether partial or complete, increases the distribution difference between the original and generated data, leading to degraded classifier performance.
Table 3: Classification performance on different training datasets.
@@ 4.2.2 Validity of Regression‑Imputed Data Versus Noise Baselines
4.2.3 Impact of Cross‑Domain Imputed Modalities on Classification Performance
@@
Table 4 summarizes the classification accuracy of LSTM, CNN1D, and MLP models under the four scenarios. The results demonstrate that our targeted imputer effectively recovers missing modalities and that each added channel contributes enriches feature extraction, thereby boosting fatigue detection performance. Importantly, this feature‑level augmentation enables leveraging knowledge derived from high‑fidelity but deployment‑constrained sensors (e.g., ECG and EEG), thereby strengthening the performance of the practical sensor configuration (i.e., VPFD) without modifying its original features.
Table 4: Classification accuracy (%) under VPFD augmentation scenarios
Dataset MLP(%) LSTM(%) CNN1D(%)
@@
VPFD w/ ECG+ECG 91.13 92.12 88.43
4.2.4 Ablation of Batch Normalization and Jacobian Regularization
The fourth experiments, summarized in Table 5, examines how different regularization strategies influence the quality of imputed modalities, ultimately, the resulting classification performance. The MLP classifier is used as the testbed in this experiment. An ablation study compares four settings—Batch Normalisation (BN) alone, Jacobian‑norm regularization alone, both combined, and neither. BN smooths the feature distribution, while the Jacobian term sharpens the network’s sensitivity to the subtle variations required for reconstructing missing channels. When applied together, these mechanisms complement each other: BN provides a stable representation space and the Jacobian constraint enhances gradient information, allowing the model to recover absent modalities most faithfully and to achieve the best overall accuracy. This complementary effect is particularly evident in the configuration where both auxiliary sources are available, as the inclusion of multiple reconstructed modalities (i.e., EEG+ECG) accentruates the performance gap bewteen models wih and without knowledge transfer. Although this study uses the MLP classifier for clarity of presentation, the underlying analysis indicates that the same synergy between distributional and gradient‑based constraints is expected to generalize to other classifiers.
Table 5: Ablation Study Results (Accuracy %)
@@ Appendix A Theoretical Analysis
Definition 2 (Ideal Classifier)
@@
Definition 3 ($\mathcal{H}\Delta\mathcal{H}$ distance)

### v2

@@ IV-B1 Regression‑Based Imputation Model Selection
IV-B2 Validity of Regression‑Imputed Data Versus Noise Baselines
The second set of experiments is designed to not only verify Theorems 1 and 2 but also demonstrate that imputed data can effectively support classifier tasks. In this case, VPFD is chosen with its gaze modality masked and reconstructed using three different methods: MLP-predicted values, MLP prediction with Gaussian noise, and pure random noise. We refer to the modified datasets as $\text{VPFD}_{\text{MLP}}$, $\text{VPFD}_{\text{MLP+noise}}$, and $\text{VPFD}_{\text{noise}}$, respectively. The added noise follows a zero‑mean Gaussian distribution whose standard deviation is twice the largest absolute value in the original VPFD gaze data. Finally, we train four separate MLP classifiers using 80% of the original VPFD dataset and each of the three modified versions, respectively. All models are evaluated on the same test set—20% of the original VPFD data—to ensure a consistent basis for comparison. Table III clearly shows that the classifier trained on regression-generated data maintains a similar performance to the one trained on the original data. This demonstrates the potential of data synthesis as a solution for addressing missing data. However, the introduction of noise, whether partial or complete, increases the distribution difference between the original and generated data, leading to degraded classifier performance.
TABLE III: Classification performance on different training datasets.
@@ IV-B2 Validity of Regression‑Imputed Data Versus Noise Baselines
IV-B3 Impact of Cross‑Domain Imputed Modalities on Classification Performance
@@
Table IV summarizes the classification accuracy of LSTM, CNN1D, and MLP models under the four scenarios. The results demonstrate that our targeted imputer effectively recovers missing modalities and that each added channel contributes enriches feature extraction, thereby boosting fatigue detection performance. Importantly, this feature‑level augmentation enables leveraging knowledge derived from high‑fidelity but deployment‑constrained sensors (e.g., ECG and EEG), thereby strengthening the performance of the practical sensor configuration (i.e., VPFD) without modifying its original features.
TABLE IV: Classification accuracy (%) under VPFD augmentation scenarios
Dataset MLP LSTM CNN1D
@@
VPFD w/ EEG+ECG 91.13 92.12 88.43
IV-B4 Ablation of Batch Normalization and Jacobian Regularization
The fourth experiments, summarized in Table V, examines how different regularization strategies influence the quality of imputed modalities, ultimately, the resulting classification performance. The MLP classifier is used as the testbed in this experiment. An ablation study compares four settings—Batch Normalisation (BN) alone, Jacobian‑norm regularization alone, both combined, and neither (serving as the baseline). BN smooths the feature distribution, while the Jacobian term sharpens the network’s sensitivity to the subtle variations required for reconstructing missing channels. When applied together, these mechanisms complement each other: BN provides a stable representation space and the Jacobian constraint enhances gradient information, allowing the model to recover absent modalities most faithfully and to achieve the best overall accuracy. This complementary effect is particularly evident in the configuration where both auxiliary sources are available, as the inclusion of multiple reconstructed modalities (i.e., EEG+ECG) accentruates the performance gap bewteen models wih and without knowledge transfer. Although this study uses the MLP classifier for clarity of presentation, the underlying analysis indicates that the same synergy between distributional and gradient‑based constraints is expected to generalize to other classifiers.
TABLE V: Ablation study results (Accuracy %)
@@ Appendix A Theoretical Analysis
Definition 2 (Ideal Classifier).
@@
Definition 3 ($\mathcal{H}\Delta\mathcal{H}$ distance).